OpenStudy (anonymous):

find the sum of the measures of exterior angles, one at each vortex, of an octagon. a. 180 degrees b. 360 degrees c. 1080 degrees d. 1440 degrees

OpenStudy (anonymous):

@legacy101

OpenStudy (anonymous):

i think it is 360 but im not sure

OpenStudy (anonymous):

I would say 180 but I am not sure :( 180 would be my best guess but I have no clue.

OpenStudy (anonymous):

how many sides does an octagon have? the sum of all interiro angles = (n-2)180 where, n equals the number of sides in the shape

OpenStudy (anonymous):

i still dont understand, can you explain a little more? im sorry im just really bad with math!

OpenStudy (anonymous):

yeah, no problem. so an octagon has 8 sides. so we will say n=8 now we use the equation: (n-2)*180 = the sum of all interior angles recal that n=8, so (8-2)*180 = ? <--- this will be the answer to your quesiton ^_^

OpenStudy (anonymous):

6x180=1080 degrees?

OpenStudy (anonymous):

good so far! "find the sum of the measures of exterior angles, one at each vortex, of an octagon. so, we just found the sum of INTERIROR angles is 1080, now we need to know what each angle is, so 1080/8 = 135. the exteriro angle at each vertex is 135 degrees, if we have 8 vertrex then, 8*135 = ?

OpenStudy (anonymous):

1080?